Hi I really need help with these questions please. Could someone show their solution too please. Thank you so much.
1.)A ball is thrown upward with an initial velocity 60m/sec. The speed of the ball decreases uniformly with time and after 3 secs its velocity is 30.6m/sec.
a.)Determine the equation relating the velocity of the ball and time
b.)Draw the line given by this equation
c.)Determine the time at which the velocity of the ball is zero

Answer by Alan3354(69443)

1.)A ball is thrown upward with an initial velocity 60m/sec. The speed of the ball decreases uniformly with time and after 3 secs its velocity is 30.6m/sec.
a.)Determine the equation relating the velocity of the ball and time
acceleration = change in speed/time
a = (30.6 - 60)/3
a = - 9.8 m/sec/sec (negative since it's decreasing)
--------------
v = 60 + a*t
v = 60 - 9.8t
========================================
b.)Draw the line given by this equation
It's a straight line thru the 2 points (0,60) and (3,30.6)
==========================================
c.)Determine the time at which the velocity of the ball is zero
v = 60 + at
v = 60 - 9.8t = 0
t = 60/9.8
t =~ 6.122 seconds